OpenStudy (anonymous):

Yves has bought 420 tropical fish for a museum display. He bought 6 times as many parrotfish as angelfish. How many of each type of fish did he buy? (Points : 1)

OpenStudy (anonymous):

if you let x = the number of angelfish then it is 6x for the number of parrotfish... and their sum is 420...

OpenStudy (anonymous):

think you can derived the needed relationship?

OpenStudy (anonymous):

\[6x+x=420\]\[7x=420\]\[x=\frac{420}{7}\]\[x=60~for~the~angelfish\]\[6x=6(60)=360~for~the~parrotfish\]to check:\[360+60=420\]

OpenStudy (anonymous):

hope this clear... :)